ICM 1998, Berlin, Aug. 18–27

Abstracts of Short Communications and Posters

Section:

8. Analysis

1991 MS Classification: 40 Seydel, Detlef, FH-Braunschweig/Wolfenbüttel, Germany Matrix Maps of Quasiconvex Null Sequences Stieglitz and Tietz identify the space q^{α} of all quasiconvex convergent sequences as a BK-space and characterize all infinite matrices mapping q^{α} into an arbitrary FK-space. We characterize all matrices mapping the space q_{0}^{α} := q^{α} ∩ c_{0} of all quasiconvex null sequences into an arbitrary FK-space . At first we consider a special matrix map of q_{0}^{α} which gives us the BKtopology on q_{0}^{α}. Further we get the β-duals of q_{0}^{α} and q^{α}. Reference: M. S tieglitz and H. Tietz, Matrixtransformationen quasikonvexer Folgen. Hokkaido Math. J. 6, 10–15 (1977).